find the value of a that makes the ordered pair a solution of the equation.

y=3x+7; (-3,a)
2x+4y=14; (-5,a)
y=11-7x; (a,-10)
9x-5y=-9; (a-1,9)

Answer by ewatrrr(24785)


Hi
Note: ordered pairs are given as (x,y)
y = 3x+7 (-3,a) |looking for y value
y = 3*-3 + 7 y = -2
2x+4y=14 (-5,a) |looking for y value
2*-5 + 4y = 14 y = 6
y = 11-7x (a,-10) |looking for x value
-10 = 11-7x
7x = 21 x = 3
9x - 5y= -9 (a-1,9) |looking for x value of ordered pair
9x - 5*9= -9
9x = 36 x = 4 | (a-1) = 4 a = 5